What is the wavelength of photons that have molar energy of 479 (kJ)/(mol)?
1
Expert's answer
2020-09-17T07:18:28-0400

Solution:

The energy of 1 photon is:

"E=\frac{E_m}{N_A}=\frac{4.79×10^5 J/mol}{6.022×10^{23} mol^{-1}}=7.95×10^{-19} J"


The energy of a photon is related to the wavelength as:

"E=\frac{h×c}{\lambda}"


Therefore:

"\lambda=\frac{h×c}{E}=\frac{6.626×10^{-34} J⋅s \,\,×\,\,3×10^8 m/s}{7.95×10^{-19} J}=2.5×10^{-7}m=250 nm"


Answer: 250 nm.